OpenStudy (kitkat16):

Help Please! A class consists of 60% men and 40% women. Blonde men compose 25% of the class, and blond women make up 20% of the class. If a student is chosen at random and is found to be a male, what is the probability that the student is blond?

OpenStudy (anonymous):

25 % , the answer is in the question otherwise i dont understand ~

OpenStudy (kitkat16):

hi kirby i do this the same way yes?

OpenStudy (kirbykirby):

Yes indeed. It's again a conditional probability.

OpenStudy (kirbykirby):

@hns109 it is not quite 25%. We know that the chosen person is male, so this is a conditional probability, so the sample space is "reduced" so we expect a higher probability.

OpenStudy (kitkat16):

25/60=5/12 = 41.67 = 0.42

OpenStudy (kirbykirby):

yep :)

OpenStudy (kitkat16):

yay!!!!!!!! Thanks for sticking around

OpenStudy (kirbykirby):

oh no problem hehe

OpenStudy (kitkat16):

:)